OpenStudy (anonymous):

Sara wants to invest $200 in an account that compounds interest monthly. She wants to have $1000 in 7 years. What should her interest rate be? i know the equation is 1000=200(1+r/12)^84 how do you solve for r

OpenStudy (anonymous):

5 = (1+r/12)^84 log 5 = 84log(1+r/12) log (1+r/12) = log 5 x 1/84 = 0.699/84 1+r/12 = 10^(0.699/84) = 1.02 r/12 = 0.02 r= 0.24